Youve decided to invest $1200 today in a local bank that is advertising a savings rate of 8% APR- guaranteed for 5 years. After 5 years, you plan to pull all of the money out of the bank, and invest with your Uncle Bob, who says hell pay you an interest rate of 6% APR for each of the next 7 years. After that, you plan to withdraw all of your money from Uncle Bob, and simply put the money under your mattress for 10 years. Hence, overall, youll be investing your money for 22 years. Part A)What interest rate are you earning during the 10-year period when your money sits under the mattress? Part B) How much money will you have after 8 years? Part C) How much money will you have after 12 years? Part D) How much money will you have after 13 years? Part E)How much money will you have after 22 years?Help with finance
Answer #1

1.
0%

2.
=1200*1.08^5*1.06^3=2099.991898

3.
=1200*1.08^5*1.06^7=2651.191388

4.
=1200*1.08^5*1.06^7=2651.191388

5.
=1200*1.08^5*1.06^7=2651.191388